# Hidden-Heavy Pentaquarks and Where to Find Them

### v1

@@ Abstract
Next, we examine B-O potentials with $(I,S)=(0,-1)$. The $\Lambda_{c}\bar{D}_{s}^{(\ast)}$ potential $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+}$ $\bm{8}$-baryon. The $\Xi_{c}\bar{D}^{(\ast)}$ potential $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+\prime}$ $\bm{8}$-baryon. The two $\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ potentials $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ must connect to the two potentials for the $\frac{3}{2}^{+}$ $\bm{8}$-baryon. The $\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ potential $\bigl{(}\frac{1}{2}^{+\prime\prime}\bigr{)}\frac{1}{2}$ must connect to the potential for an orbitally-excited $\bm{8}$-baryon, since Table 2 does not have a third $\bm{8}$-baryon with $j^{\pi}=\frac{1}{2}^{+}$ and $(I,S)=(0,-1)$. The lack of evidence for $c\bar{c}s$ pentaquarks near the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{\ast}$ thresholds suggests that the $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ potential does not support bound states. The two observed $c\bar{c}s$ pentaquarks are near the $\Xi_{c}\bar{D}$ and $\Xi_{c}\bar{D}^{\ast}$ thresholds. This suggests that the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential supports bound states.
@@
The first-order correction in $1/m_{Q}$ to the B-O potentials depends on $\bm{r}$ and the spin vectors $\bm{S}_{Q}$ and $\bm{S}_{\bar{Q}}$ of the heavy quark and antiquark. It includes short-distance terms that give spin splittings to the $\bm{8}$-baryons, intermediate-distance terms that include spin-orbit, spin-spin, and tensor interactions, and long-distance terms that give spin splittings to well-separated heavy hadrons. At large distances, the correction is a spin-splitting potential $V_{\text{SS}}$ that does not depend on $\bm{r}$. For the pentaquark B-O potentials in Table 1,
@@
where $\bm{j}_{\text{B}}$ and $\bm{j}_{\text{M}}$ are the spins of the $\bm{3}$-baryon and $\bm{\bar{3}}$-meson and $\Delta_{\text{B}}$ and $\Delta_{\text{M}}$ are the spin splittings of the heavy baryons and heavy mesons.
A simple approximation to the spin splittings of the pentaquarks can be obtained using first-order perturbation theory in $V_{\text{SS}}$. The spin splittings are calculated by diagonalizing the matrix $V_{\text{SS}}$ for the states in the HQSS multiplet. We set the orbital angular momentum of the $Q\bar{Q}$ pair to zero since the ground state in the B-O potentials is expected to be $S$-wave.
We identify the four observed $c\bar{c}$ pentaquarks as the ground-state quartet for the $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ potentials that approach the $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ threshold at large $r$. We predict the $J^{P}$ quantum numbers for $P_{c\bar{c}}(4312)^{+}$, $P_{c\bar{c}}(4380)^{+}$, $P_{c\bar{c}}(4440)^{+}$, and $P_{c\bar{c}}(4457)^{+}$ to be $\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{3}{2}^{-}$, and $\frac{5}{2}^{-}$, respectively. We identify the energy of the multiplet in the absence of spin splittings with the spin-weighted average of the central values of the measured masses, which we calculate using their predicted quantum numbers. We approximate the spin splittings by first-order perturbation theory in $V_{\text{SS}}$ with $\Delta_{\text{B}}=66.5$ MeV and $\Delta_{\text{M}}=142.1$ MeV. The energy level diagram is illustrated in Fig. 2.
@@
We identify the two observed $c\bar{c}s$ pentaquarks as members of the ground-state triplet for the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential that approaches the $\Xi_{c}\bar{D}^{(\ast)}$ threshold at large $r$. The third pentaquark is predicted to have a mass close to that of $P_{c\bar{c}s}(4459)^{0}$. The quantum numbers of $P_{c\bar{c}s}(4338)^{0}$ are $\frac{1}{2}^{-}$, in agreement with experiment [18]. We predict the quantum numbers of $P_{c\bar{c}s}(4459)^{0}$ and the third pentaquark to be $\frac{1}{2}^{-}$ and $\frac{3}{2}^{-}$ or $\frac{3}{2}^{-}$ and $\frac{1}{2}^{-}$. We identify the energy of the multiplet in the absence of spin splittings with the spin-weighted average of the central values of the measured masses, which we calculate using the predicted quantum numbers for $P_{c\bar{c}s}(4459)^{0}$ and the third pentaquark with the same mass. We approximate the spin splittings by first-order perturbation theory in $V_{\text{SS}}$ with $\Delta_{\text{B}}=0$ and $\Delta_{\text{M}}=142.1$ MeV. The energy level diagram is illustrated in Fig. 3.
@@
Figure 3: Energy levels for the ground-state triplet of $c\bar{c}s$ pentaquarks in the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential with $(I,S)=(0,-1)$ (thick bars) versus the measured masses (shaded boxes with thin lines).
@@
Pentaquark Decays. The decays of hidden-heavy pentaquarks into heavy-hadron pairs could produce large partial widths unless they are suppressed. The observed pentaquarks, except for $P_{c\bar{c}}(4380)^{+}$, have surprisingly narrow decay widths. It is therefore important to identify a suppression mechanism for kinematically allowed decays into charm-hadron pairs for each narrow pentaquark. The decay into a constituent heavy-hadron pair can be calculated using first-order perturbation theory in $V_{\text{SS}}$. The decay rate is the product of the probability for that heavy-hadron pair as a constituent, the square of the expectation value of $V_{\text{SS}}$, and a phase space factor. The decay into a lower-energy heavy-hadron pair can proceed through a transition between B-O potentials. Model-independent relations for these decays can be obtained using the techniques introduced in Ref. [19].
@@
The $c\bar{c}$ pentaquarks can decay into $\Lambda_{c}\bar{D}$ and $\Lambda_{c}\bar{D}^{\ast}$ through a transition between the $\bigl{(}\frac{3}{2}^{+}\bigr{)}\frac{1}{2}$ and $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ potentials in Table 1. The relations of Ref. [19] indicate that these decays have $D$-wave suppression, which is a remarkable prediction of the B-O framework. $P_{c\bar{c}}(4380)^{+}$, $P_{c\bar{c}}(4440)^{+}$, and $P_{c\bar{c}}(4457)^{+}$ are also kinematically allowed to decay into the constituents $\Sigma_{c}\bar{D}$ and $\Sigma_{c}^{\ast}\bar{D}$. The only decays that can proceed through an $S$-wave channel are $P_{c\bar{c}}(4380)^{+}$ and $P_{c\bar{c}}(4440)^{+}$ into $\Sigma_{c}^{\ast}\bar{D}$. The mass of $P_{c\bar{c}}(4380)^{+}$ is below the $\Sigma_{c}^{\ast}\bar{D}$ threshold, but it can decay into $\Sigma_{c}^{\ast}\bar{D}$ because it has a large width and much of its line shape is above the threshold. The probabilities for $\Sigma_{c}^{\ast}\bar{D}$ in $P_{c\bar{c}}(4380)^{+}$ and $P_{c\bar{c}}(4440)^{+}$ are 0.611 and 0.056. The expectation values of $V_{\text{SS}}$ are equal. Thus the ratio of the decay rates is the ratio of the phase-space factors multiplied by 10.9. The ratio of the measured widths of $P_{c\bar{c}}(4380)^{+}$ and $P_{c\bar{c}}(4440)^{+}$ is $10.0^{+6.4}_{-6.9}$. The compatibility with this large ratio comes primarily from the small probability 0.056 for the $\Sigma_{c}^{\ast}\bar{D}$ component of $P_{c\bar{c}}(4440)^{+}$, which is a remarkable prediction of the B-O framework.
@@
$P_{c\bar{c}s}(4459)^{0}$ is also kinematically allowed to decay into the constituents $\Xi_{c}\bar{D}$, but the probability for the $\Xi_{c}\bar{D}$ component is zero indicating $D$-wave suppression. The ratio of the measured widths of $P_{c\bar{c}s}(4459)^{0}$ and $P_{c\bar{c}s}(4338)^{0}$ is $2.5^{+1.6}_{-1.4}$, which is compatible with the ratio in Eq. (2). The transition amplitude between the $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ and $\bigl{(}\frac{1}{2}^{+}\bigr{)}\frac{1}{2}$ potentials is proportional to their flavor overlap. The flavors of $\Xi_{c}\bar{D}^{(\ast)}$ and $\Lambda_{c}\bar{D}_{s}^{(\ast)}$ are orthogonal, so the transition amplitude approaches zero at large $r$. The transition amplitude at small $r$ is determined by the mixing between the SU(3) singlet and the $(I,S)=(0,-1)$ member of the SU(3) octet of $\bm{8}$-baryons with $j^{\pi}=\frac{1}{2}^{+}$. The small width for the $c\bar{c}s$ pentaquarks into $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{\ast}$ could be explained by the decays being dominated by a region of $r$ where the flavor overlap is small.
Molecular Picture. The molecular picture corresponds in the B-O framework to the assumption that all the B-O potentials that approach the same static-hadron-pair threshold at large $r$ support bound states. If these potentials are the same as those associated with a single low-energy $\bm{8}$-baryon at small $r$, the B-O approximation will be qualitatively similar to the molecular picture. An example is $c\bar{c}s$ pentaquarks with $(I,S)=(0,-1)$. If there are additional B-O potentials that approach the static-hadron-pair threshold at large $r$, the B-O approximation may be qualitatively different from the molecular picture. An example is $c\bar{c}$ pentaquarks with $(I,S)=\bigl{(}\frac{1}{2},0\bigr{)}$. The molecular picture predicts seven ground-state $c\bar{c}$ pentaquarks compared to the B-O prediction of four.
@@
Summary. We have applied the pattern of exotic hidden-heavy hadrons proposed in Ref. [13] to hidden-charm pentaquarks. The four observed $c\bar{c}$ pentaquarks form the HQSS quartet for the $c\bar{c}$ ground state in $\bigl{(}\frac{3}{2}^{+}\bigr{)}\Lambda$ potentials. The two observed $c\bar{c}s$ pentaquarks belong to the HQSS triplet for the $c\bar{c}$ ground state in a $\bigl{(}\frac{1}{2}^{+\prime}\bigr{)}\frac{1}{2}$ potential. Our prediction that there are only four $c\bar{c}$ pentaquarks in the ground-state multiplet differs from most previous studies. Our most surprising prediction is that $P_{c\bar{c}}(4457)^{+}$ has quantum numbers $\frac{5}{2}^{-}$. To the best of our knowledge, all previous predictions in the literature for the quantum numbers of $P_{c\bar{c}}(4457)^{+}$ have been $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$. We also explained why $P_{c\bar{c}}(4380)^{+}$ is the only observed pentaquark with a large total decay width.

### v3

@@ Abstract
Next, we examine B-O potentials with (I,S)=(0,1)(I,S)=(0,-1). The $\Lambda_{c}\bar{D}_{s}^{(\ast)}$ potential $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+}$ $\bm{8}$-baryon. The $\Xi_{c}\bar{D}^{(\ast)}$ potential $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ must connect to the potential for the $\frac{1}{2}^{+\prime}$ $\bm{8}$-baryon. The $\frac{1}{2}^{+}$ and $\frac{1}{2}^{+\prime}$ $\bm{8}$-baryons are orthogonal linear combinations of the ground-state $\bm{8}$-baryons in the SU(3) singlet and the SU(3) octet. The two $\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ potentials $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ must connect to the two potentials for the $\frac{3}{2}^{+}$ $\bm{8}$-baryon. The $\Xi_{c}^{\prime(\ast)}\bar{D}^{(\ast)}$ potential $\bigl(\frac{1}{2}^{+\prime\prime}\bigr)\frac{1}{2}$ must connect to the potential for an orbitally excited $\bm{8}$-baryon, since Table 2 does not have a third $\bm{8}$-baryon with $j^{\pi}=\frac{1}{2}^{+}$ and (I,S)=(0,1)(I,S)=(0,-1). The lack of evidence for $c\bar{c}s$ pentaquarks near the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{\ast}$ thresholds suggests that the $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ potential does not support bound states. The two observed $c\bar{c}s$ pentaquarks are near the $\Xi_{c}\bar{D}$ and $\Xi_{c}\bar{D}^{\ast}$ thresholds. This suggests that the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential supports bound states.
@@
The first-order correction in $1/m_{Q}$ to the B-O potentials depends on $\bm{r}$ and the spin vectors $\bm{S}_{Q}$ and $\bm{S}_{\bar{Q}}$ of the heavy quark and antiquark. It includes short-distance terms that give spin splittings to the $\bm{8}$-baryons and long-distance terms that give spin splittings to well-separated heavy hadrons. At large distances, the correction is a spin-splitting potential $V_{\text{SS}}$ that does not depend on $\bm{r}$. For the pentaquark B-O potentials in Table 1,
@@
where $\bm{j}_{\text{B}}$ and $\bm{j}_{\text{M}}$ are the spins of the $\bm{3}$-baryon and $\bm{\bar{3}}$-meson and $\Delta_{\text{B}}$ and $\Delta_{\text{M}}$ are the spin splittings of the heavy baryons and heavy mesons. A simple approximation to the spin splittings of the pentaquarks can be obtained using first-order perturbation theory in $V_{\text{SS}}$. The spin splittings are calculated by diagonalizing the matrix $V_{\text{SS}}$ for the states in the HQSS multiplet. We set the orbital angular momentum of the $Q\bar{Q}$ pair to zero since the ground state in the B-O potentials is expected to be mostly $S$-wave.
We identify the three observed $c\bar{c}$ pentaquarks as members of the ground-state quartet in the $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ potentials that approach the $\Sigma_{c}^{(\ast)}\bar{D}^{(\ast)}$ threshold at large $r$. We predict the $J^{P}$ quantum numbers for $P_{c\bar{c}}(4312)^{+}$, $P_{c\bar{c}}(4440)^{+}$, and $P_{c\bar{c}}(4457)^{+}$ to be $\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, and $\frac{5}{2}^{-}$, respectively. We also predict an additional $\frac{3}{2}^{-}$ state. We identify the energy of the multiplet in the absence of spin splittings with the spin-weighted average of the central values of the measured masses, which we calculate using their predicted quantum numbers. We approximate the spin splittings by first-order perturbation theory in $V_{\text{SS}}$. We determine $\Delta_{\text{B}}=66.5$ MeV from the mass splitting between $\Sigma_{c}^{\ast}$ and $\Sigma_{c}$ and $\Delta_{\text{M}}=142.1$ MeV from the mass splitting between $\bar{D}^{\ast}$ and $\bar{D}$. The energy level diagram is illustrated in Figure 3. The additional $\frac{3}{2}^{-}$ state is predicted to be near the $\Sigma_{c}^{\ast}\bar{D}$ threshold. We denote it by $P_{c\bar{c}}(E)^{+}$ with $E$ near 4385.
@@
We identify the two observed $c\bar{c}s$ pentaquarks as members of the ground-state triplet for the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential that approaches the $\Xi_{c}\bar{D}^{(\ast)}$ threshold at large $r$. The third pentaquark is predicted to have a mass close to that of $P_{c\bar{c}s}(4459)^{0}$. The quantum numbers of $P_{c\bar{c}s}(4338)^{0}$ are $\frac{1}{2}^{-}$, in agreement with experiment [19]. We predict the quantum numbers of $P_{c\bar{c}s}(4459)^{0}$ and the third pentaquark to be $\frac{1}{2}^{-}$ and $\frac{3}{2}^{-}$ or $\frac{3}{2}^{-}$ and $\frac{1}{2}^{-}$. We identify the energy of the multiplet in the absence of spin splittings with the spin-weighted average of the central values of the measured masses, which we calculate using the predicted quantum numbers for $P_{c\bar{c}s}(4459)^{0}$ and the third pentaquark with the same mass. We approximate the spin splittings by first-order perturbation theory in $V_{\text{SS}}$. We determine $\Delta_{\text{M}}=142.1$ MeV from the mass splitting between $\bar{D}^{\ast}$ and $\bar{D}$ and we set $\Delta_{\text{B}}=0$ since $\Xi_{c}$ is a HQSS singlet. The energy level diagram is illustrated in Figure 4.
@@
Figure 4: Energy levels for the ground-state triplet of $c\bar{c}s$ pentaquarks in the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential with (I,S)=(0,1)(I,S)=(0,-1) (thick bars) versus the measured masses (shaded boxes with thin lines). The predicted quantum numbers of $P_{c\bar{c}s}(4459)^{0}$ are either $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$.
@@
Pentaquark Decays. The decays of hidden-heavy pentaquarks into heavy-hadron pairs could produce large partial widths unless they are suppressed. The observed pentaquarks have surprisingly narrow decay widths. It is therefore important to identify a suppression mechanism for the kinematically allowed decays into charm-hadron pairs for each narrow pentaquark. The decay into a constituent heavy-hadron pair can be calculated using first-order perturbation theory in $V_{\text{SS}}$. The decay rate is the product of the probability for that heavy-hadron pair as a constituent, the square of the matrix element of $V_{\text{SS}}$, and a phase space factor. The decay into a lower-energy heavy-hadron pair can proceed through a transition between B-O potentials. Model-independent relations for these decays can be obtained using the techniques introduced in Ref. [21].
@@
The $c\bar{c}$ pentaquarks can decay into $\Lambda_{c}\bar{D}$ and $\Lambda_{c}\bar{D}^{\ast}$ through a transition between the $\bigl(\frac{3}{2}^{+}\bigr)\frac{1}{2}$ and $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ potentials in Table 1. The relations of Ref. [21] indicate that these decays have $D$-wave suppression, which is a remarkable prediction of the B-O framework. $P_{c\bar{c}}(4440)^{+}$ and $P_{c\bar{c}}(4457)^{+}$ are kinematically allowed to decay into the constituents $\Sigma_{c}\bar{D}$ and $\Sigma_{c}^{\ast}\bar{D}$. The fourth $c\bar{c}$ pentaquark $P_{c\bar{c}}(E)^{+}$ is kinematically allowed to decay into $\Sigma_{c}\bar{D}$ and also into $\Sigma_{c}^{\ast}\bar{D}$ if $E>4385$. The only decays that can proceed through an $S$-wave channel are $P_{c\bar{c}}(E)^{+}$ and $P_{c\bar{c}}(4440)^{+}$ into $\Sigma_{c}^{\ast}\bar{D}$. The probabilities for $\Sigma_{c}^{\ast}\bar{D}$ in $P_{c\bar{c}}(E)^{+}$ and $P_{c\bar{c}}(4440)^{+}$, which are determined by $\Delta_{\text{B}}/\Delta_{\text{M}}$, are 0.611 and 0.056. The tiny $\Sigma_{c}^{\ast}\bar{D}$ probability for $P_{c\bar{c}}(4440)^{+}$ is a remarkable prediction of the B-O framework. The matrix elements of $V_{\text{SS}}$ factor out of the ratio. If $E>4385$, the ratio of the decay rates is the ratio of the phase-space factors multiplied by the probability ratio 10.9.
@@
This prediction is just the phase-space ratio and its uncertainty comes from the pentaquark masses. $P_{c\bar{c}s}(4459)^{0}$ is also kinematically allowed to decay into the constituents $\Xi_{c}\bar{D}$, but the probability for the $\Xi_{c}\bar{D}$ component is zero indicating $D$-wave suppression. The ratio of the measured widths of $P_{c\bar{c}s}(4459)^{0}$ and $P_{c\bar{c}s}(4338)^{0}$ is $2.5^{+1.6}_{-1.4}$, which is compatible with the ratio in Equation (2). The transition amplitude between the $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ and $\bigl(\frac{1}{2}^{+}\bigr)\frac{1}{2}$ potentials is proportional to their flavor overlap. The flavors of $\Xi_{c}\bar{D}^{(\ast)}$ and $\Lambda_{c}\bar{D}_{s}^{(\ast)}$ are orthogonal, so the transition amplitude approaches zero at large $r$. The transition amplitude at small $r$ is determined by the mixing between the SU(3) singlet and the (0,1)(0,-1) member of the SU(3) octet of $\bm{8}$-baryons with $j^{\pi}=\frac{1}{2}^{+}$. The small width for the $c\bar{c}s$ pentaquarks into $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{\ast}$ could be explained by the decays being dominated by a region of $r$ where the flavor overlap is small.
Molecular Picture. The molecular picture corresponds in the B-O framework to the assumption that all the B-O potentials that approach the same static-hadron-pair threshold at large $r$ support bound states. If these potentials are the same as those associated with a single low-energy $\bm{8}$-baryon at small $r$, the B-O approximation will be qualitatively similar to the molecular picture. An example is $c\bar{c}s$ pentaquarks with (I,S)=(0,1)(I,S)=(0,-1). If there are additional B-O potentials that approach the static-hadron-pair threshold at large $r$, the B-O approximation may be qualitatively different from the molecular picture. An example is $c\bar{c}$ pentaquarks with $(I,S)=\bigl(\frac{1}{2},0\bigr)$. The molecular picture predicts seven ground-state $c\bar{c}$ pentaquarks compared to the B-O prediction of four. A recent study has considered an alternative B-O approximation that reproduces the molecular picture by ignoring the requirement that the $\frac{1}{2}^{+\prime}$ potential must connect to that for an orbitally excited $\bm{8}$-baryon [22].
@@
Summary. We have applied the pattern of exotic hidden-heavy hadrons proposed in Ref. [14] to hidden-heavy pentaquarks. We have used the B-O symmetries to connect hadron-pair potentials at long distances with $\bm{8}$-hadron potentials at short distances. We have used these connections together with experimental evidence to infer which B-O potentials may support bound states. The three observed $c\bar{c}$ pentaquarks belong to the HQSS quartet for the $c\bar{c}$ ground state in $\bigl(\frac{3}{2}^{+}\bigr)\Lambda$ potentials. The two observed $c\bar{c}s$ pentaquarks belong to the HQSS triplet for the $c\bar{c}$ ground state in a $\bigl(\frac{1}{2}^{+\prime}\bigr)\frac{1}{2}$ potential. Our most surprising prediction is that $P_{c\bar{c}}(4457)^{+}$ has quantum numbers $\frac{5}{2}^{-}$, which agrees with only a few previous predictions [23, 24, 25]. Most previous studies have predicted $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$; see, for instance, Refs. [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] and references therein. Our prediction could be verified by resolving the angular distribution of the decay products in the discovery channel $P_{c\bar{c}}(4457)^{+}\to J/\psi\,p$. If the quantum numbers of $P_{c\bar{c}}(4457)^{+}$ are $\frac{1}{2}^{-}$ or $\frac{3}{2}^{-}$, the decay can proceed in $S$-wave. Our assignment $\frac{5}{2}^{-}$ rules out $S$-wave, and we instead predict that the decay $P_{c\bar{c}}(4457)^{+}\to J/\psi\,p$ will be dominated by $D$-wave.